OpenStudy (anonymous):

Use the quadratic formula to solve the problem. A soccer player throws the ball onto the field from the sideline. An equation modeling the ball's flight is h = -16t2 + 18t + 2. In how many seconds will the ball reach the ground? A. in about 2 seconds B. in about 1.2 seconds C. in about 1.8 seconds D. Cannot be determined

Directrix (directrix):

The ball hits the ground when h = 0. The question is to find when -16t2 + 18t + 2= 0. -16t2 + 18t + 2= 0 Divide through by -2 8t^2 - 9t - 1 = 0 This quadratic appears not to factor. So, use the quadratic formula to find the roots, one of which will be the answer to the question.

OpenStudy (anonymous):

^_v huh

Directrix (directrix):

Read about the quadratic formula at the link. http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

I already did but I don't understand it

Directrix (directrix):

Well, copy the quadratic formula and post it here. Then, we'll figure out the a, b, and c values. If you do not want to use the quadratic formula, then suggest some other technique and I'll try to help you with that.

OpenStudy (anonymous):

ok